OpenStudy (aravindg):

does a=b always imply 1/a=1/b?

OpenStudy (anonymous):

yes,

OpenStudy (blockcolder):

Except of course when a=b=0 but otherwise, yeah.

OpenStudy (aravindg):

@badreferences , @apoorvk , @dumbcow , @dpaInc

OpenStudy (aravindg):

hmm...

OpenStudy (blockcolder):

If a=b, and they're not zero, you can divide both sides by ab to get 1/a=1/b.

OpenStudy (aravindg):

@EarthCitizen , @Ishaan94 , @Mani_Jha

OpenStudy (aravindg):

oh!! thats the answer i was looking for!!@blockcolder

OpenStudy (aravindg):

@blockcolder

OpenStudy (aravindg):

this seems a simple qn ..bt i had this doubt from my small classes

OpenStudy (earthcitizen):

a=b, so long as a \[ a \neq0\]

OpenStudy (anonymous):

It's been answered, but a more rigorous way of answering it would be:\[\left\{a=b\mid\forall \left(ab\neq0\right)\right\}\]

OpenStudy (anonymous):

Whoops, I mean:\[\left\{a=b\therefore\frac{1}{a}=\frac{1}{b}\mid\forall \left(ab\neq0\right)\right\}\]

OpenStudy (anonymous):

As either \(a,b\) can be \(0\), but not necessarily both, for the implication to be demonstrably false. A simpler way of showing this is by determining that the product must not be zero.